# Jr. AI Scientist and Its Risk Report: Autonomous Scientific Exploration from a Baseline Paper

### v2

@@ Abstract
Understanding the current capabilities and risks of AI Scientist systems is essential for ensuring trustworthy and sustainable AI-driven scientific progress while preserving the integrity of the academic ecosystem. To this end, we develop Jr. AI Scientist, a state-of-the-art autonomous AI scientist system that mimics the core research workflow of a novice student researcher: Given the baseline paper from the human mentor, it analyzes its limitations, formulates novel hypotheses for improvement, and iteratively conducts experiments until improvements are realized, and writes a paper with the results. Unlike previous approaches that assume full automation or operate on small-scale code, Jr. AI Scientist follows a well-defined research workflow and leverages modern coding agents to handle complex, multi-file implementations, leading to scientifically valuable contributions. Through our experiments, the Jr. AI Scientist successfully generated new research papers that build upon real NeurIPS, IJCV, and ICLR works by proposing and implementing novel methods. For evaluation, we conducted automated assessments using AI Reviewers, author-led evaluations, and submissions to Agents4Science, a venue dedicated to AI-driven scientific contributions. The findings demonstrate that Jr. AI Scientist generates papers receiving higher review scores than existing fully automated systems. Nevertheless, we identify important limitations from both the author evaluation and the Agents4Science reviews, indicating the potential risks of directly applying current AI Scientist systems and key challenges for future research. Finally, we comprehensively report various risks identified during development. We believe this study clarifies the current role and limitations of AI Scientist systems, offering insights into the areas that still require human expertise and the risks that may emerge as these systems evolve.
Note: This is a comprehensive report on our AI Scientist project. We do not recommend the use of the AI Scientist for generating conference submission papers. Our primary objective is to share insights gained from both successful cases and observed risks with the community to foster a deeper understanding of AI Scientists.
@@ 1 Introduction
Understanding the current upper bound of capabilities in AI Scientist systems, autonomous agents capable of conducting scientific research, is crucial for promoting sustainable, AI-driven scientific progress. Nevertheless, developers must remain conscious of the potential risks these systems pose to the academic ecosystem and commit to advancing them responsibly. Since 2025, a new venue dedicated to evaluating AI-driven scientific contributions, the Agents4Science conference (agents4science), has emerged. Through such a platform, developers of AI Scientist systems are encouraged to engage in responsible research and development, ensuring both the protection of the academic ecosystem and the long-term sustainability of scientific progress.
@@
In recent years, several works have explored the concept of AI Scientists (lu2024ai, weng2024cycleresearcher, yamada2025ai, tang2025ai). However, the quality of research papers produced by these systems remains insufficient. One major reason is that the problem setting of achieving fully automated science is overly ambitious and often lacks clearly defined scientific goals for AI Scientists. Without a specific goal, these systems tend to generate undirected discoveries that appear to lack genuine scientific value. Another limitation is that current systems are limited to small-scale code experiments (zhu2025ai, lu2024ai, yamada2025ai), lacking the scale and complexity needed for meaningful science. Achieving real scientific contributions requires not just ideas but strong implementation capability to handle complex codebases.
@@
As an initial step toward enabling AI Scientists to produce genuine scientific value, we can take inspiration from how student researchers begin their research. When a student first joins a research lab, a common and meaningful process is as follows: the mentor assigns a key paper, the student analyzes its limitations, proposes an improvement hypothesis, implements the idea on the baseline code, iteratively conducts experiments until improvements are realized, and finally writes a paper summarizing the results. Through this process, the student learns the fundamental workflow of scientific research and gains the skills and experience needed for more creative work later on. Also, improving a baseline method is not only an important stage in early research training but also a valuable research goal in many fields where advancing task performance remains a central scientific pursuit.
@@
In this paper, we introduce Jr. AI Scientist, a new AI Scientist that mimics the essential research workflow of a novice student researcher: Starting from a baseline paper, it identifies key limitations, proposes an improvement hypothesis, validates it through rigorous experimentation, and writes a paper with the results. Table˜1 shows a comparison of the problem setting with existing research. This problem setting reframes previously ambitious goals into a more specific objective, providing a clear optimization direction for the AI Scientist. Moreover, because the framework operates on the actual codebase of baseline papers, it can generate results with genuine scientific value. These aspects collectively represent an essential first step toward the autonomous generation of reliable and high-quality research papers.
@@
Our Jr. AI Scientist consists of three main components: (1) automatic idea generation based on the limitations of a given paper, (2) automatic implementation and thorough validation of the proposed ideas, and (3) automatic writing of a research paper based on the obtained results. This system is built upon AI Scientist v2 (yamada2025ai), but our work differs from prior studies (lu2024ai, weng2024cycleresearcher, yamada2025ai, tang2025ai) in several key aspects: First, by leveraging the latest coding agents (e.g. Claude Code (claude_code)), our system can handle realistic multi-file codebases, which were difficult to process in previous AI Scientist Systems. Second, by incorporating the full set of resources from a given baseline paper, the system exploits all available artifacts such as LaTeX sources, PDFs, and codebases, thereby substantially improving the scope and quality of every stage in the research pipeline. Finally, by refining every stage of the process, our framework enables the autonomous generation of research papers that are both higher in quality and more trustworthy.
@@
AI Scientist Systems Starting Point Code Complexity Review Score AI Scientist-v1 (lu2024ai) Template code Single file 3.30 AI Scientist-v2 (yamada2025ai) General idea Single file 2.75 AI Researcher (tang2025ai) 15-20 existing papers Multiple files 3.25 Jr. AI Scientist (ours) One baseline paper and code Multiple files 5.75
@@
For the baseline papers, we selected papers for which we obtained permission from the original authors. Specifically, we used three papers: NeurIPS 2023 paper (miyai2023locoop) and IJCV 2025 paper (miyai2025gl) on out-of-distribution (OOD) detection (msp17iclr, yang2024generalized), and an ICLR 2025 spotlight paper (zhang2024min) on pre-training data detection for large language models (LLMs). Refer to Section˜4.1 for the detailed rationale behind the selection of these papers. For the evaluation, we conducted three evaluations: (1) an automated assessment using DeepReviewer (zhu2025deepreview), (2) an author-led evaluation, and (3) submission to the Agents4Science conference (agents4science). DeepReviewer automatically compared our generated papers with existing AI-generated works to assess overall quality. The author evaluation examined the outputs for hallucinations or fabricated content. Finally, the Agents4Science (agents4science) platform provides rigorous evaluation and feedback from the community platform.
@@
Through our experiments, the Jr. AI Scientist successfully generated new research papers that build upon the above top-tier venues works by proposing and implementing novel algorithms. As for the evaluation using DeepReviewer (zhu2025deepreview), the papers generated by Jr. AI Scientist achieved substantially higher review scores compared to the existing AI-generated papers. Therefore, our Jr. AI Scientist can be regarded as the most capable autonomous AI Scientist. However, we also observed that Jr. AI Scientist still exhibits some failures and unresolved challenges through the author evaluation and Agents4Science conference. To share these challenges and lessons with the research community, we analyze the feedback and evaluation results from the Agents4Science conference and include the author evaluation, which helps clarify what is required to further improve the quality of Jr. AI Scientist systems.
@@
Revealing Strengths and Limitations of Jr. AI Scientist: We conducted extensive evaluations using open-source AI reviewers, Agent4Science, and author evaluation. The results demonstrate that Jr. AI Scientist generates higher-quality research papers than existing AI Scientists, while also revealing key challenges for future improvement.
@@ 2 Related Work
Automated Scientific Discovery. Recent progress has significantly reshaped the role of AI in automating end-to-end scientific research. AI Scientist-v1 (lu2024ai) was an early milestone, showcasing how advanced language models can autonomously generate research ideas, run experiments, and draft scientific papers. This work was followed by a series of subsequent studies in machine learning fields (zochi2025, tang2025ai) and diverse scientific disciplines (villaescusa2025denario, mitchener2025kosmos) that further advanced this line of research. However, these approaches often suffer from an overly ambitious problem setting that aims to achieve fully automated science and tend to lack clearly defined scientific objectives for AI Scientists. Without specific goals, such systems often produce undirected discoveries that lack genuine scientific value. To address this issue, our Jr. AI Scientist builds on existing baselines and conducts research within a well-defined research workflow, aiming to generate higher-quality scientific papers. As concurrent work, DeepScientist (weng2025deepscientist) also adopts a baseline-based approach. DeepScientist (weng2025deepscientist) focuses mainly on experimental performance, formalizing discovery as a Bayesian Optimization problem. However, its overall framework design and workflow integration are outlined at a high level, with limited discussion of implementation details. In contrast, our Jr. AI Scientist explicitly articulates each stage of the research process and further aims to contribute to the community by comprehensively reporting the failures and risks encountered throughout scientific exploration.
@@
AI-Assisted Scientific Research. Research specialized for each element of the research process has also been actively explored (chen2025ai4research). For the idea generation phase, si2024can investigates the novelty of the LLM-generated ideas, and si2025ideation investigates the ideation–execution gap. For the survey phase, OpenScholar (asai2024openscholar) have been developed to support literature review. For the experimental phase, AlphaEvolve (novikov2025alphaevolve) leverages large-scale trial-and-error strategies to enhance the performance. For the writing and review phase, CycleResearcher (weng2024cycleresearcher) provides a learning framework specialized for scientific writing, while DeepReviewer (zhu2025deepreview) focuses on the review process. Recent studies provide a comprehensive overview of automated review, outlining key challenges, proposing a practical review pipeline for real-world implementation, and constructing a large-scale dataset to support automated review research (ZHUANG2025103332, Lin2023ASPR, lin2023moprd). Beyond these, rather than pursuing full automation like AI Scientists, AI Co-Scientist (gottweis2025towards) emphasizes collaboration between humans and AI. In this work, instead of focusing on individual parts of the research process, we investigate the entire end-to-end research cycle, aiming to rigorously evaluate both the performance and the limitations.
@@
Failures and Risk Analysis for AI Scientist Systems. There are very few studies that thoroughly analyze or report the risks and failure cases of AI Scientist systems. Although tang2024prioritizing summarizes the risks that AI Scientists may pose, it focuses on hypothesis-based potential risks, rather than empirically observed risks or failures. While beel2025evaluating provides an in-depth analysis of failure cases in AI Scientist-v1 (lu2024ai), these findings are based on the early AI Scientist, and the analysis was not conducted from the developer’s perspective. While luo2025more examines four failure modes (benchmark selection, data leakage, metric misuse, and post-hoc bias), their analysis is limited to experimental diagnostics and early AI Scientists without modern coding. Thus, their analysis remains somewhat limited in scope.
@@ 3.1 Preparation: Baseline Paper Selection
The preparation stage involves selecting a baseline paper, obtaining its LaTeX source files and PDF, and the baseline code. This setup is realistic because many recent publications are released on arXiv with LaTeX sources, and their implementation code is shared on GitHub. While there might be some augment that AI agents should automatically select a baseline and reproduce the code, current reproducibility rates from papers are still limited (siegel2024core, xiang2025scireplicate, starace2025paperbench), making complete automation premature. Since our goal is to emulate how a human scientist engages in early-stage research under the guidance of a mentor, we explicitly include this preparatory stage.
When constructing the baseline code, we followed AI Scientist v1 (lu2024ai) and made only minor modifications to the existing implementation so that the experiments could be executed via baseline.py and the results could be visualized via plot.py. Defining such an experimental entry point facilitated easier management and reproducibility of the execution process in the experimental section.
@@ 3.2 Idea Generation Phase
We provide a baseline paper as text to an LLM (e.g. GPT-o4-mini (o3-o4-mini)) and prompt it to output the limitations of the work. Based on both the baseline paper and the limitations, the LLM is then guided to propose potential research ideas. Following AI Scientist v2 (yamada2025ai), the system evaluates the originality of proposed ideas through literature review tools such as Semantic Scholar, which review papers citing the baseline work and papers with similar concepts. If conceptually similar ideas are identified, they are refined; otherwise, they are clearly distinguished from prior work. These steps define the preliminary research idea.
@@ 3.3.1 Preliminary: Coding Agent Usage
A powerful coding agent (e.g. Claude Code (claude_code)) is employed to translate research ideas into concrete implementations. We provide the coding agent with a working directory that contains the baseline implementations (prepared at Section˜3.1) and give it detailed instructions through input prompts. The agent is informed of how to use the main scripts—baseline.py, which serves as the experimental entry point, and plot.py, which visualizes the experimental results. We use claude-sonnet-4-20250514 within Claude Code (version 1.0.24).

### v4

@@ Abstract
Understanding the current capabilities and risks of AI Scientist systems (autoresearch) is essential for ensuring trustworthy and sustainable AI-driven scientific progress while preserving the integrity of the academic ecosystem. To this end, we develop Jr. AI Scientist, a state-of-the-art autonomous AI scientist system that mimics the core research workflow of a novice student researcher: Given the baseline paper from the human mentor, it analyzes its limitations, formulates novel hypotheses for improvement, iteratively experiments until improvements are achieved, and writes a paper with the results. Unlike previous approaches that assume full automation or operate on small-scale code, Jr. AI Scientist follows a well-defined research workflow and leverages modern coding agents to handle complex, multi-file implementations, leading to scientifically valuable contributions. Through our experiments, the Jr. AI Scientist successfully generated new research papers that build upon real NeurIPS, IJCV, and ICLR works by proposing and implementing novel methods. For evaluation, we conducted automated assessments using AI Reviewers, author-led evaluations, and submissions to Agents4Science, a venue dedicated to AI-driven contributions. The findings demonstrate that Jr. AI Scientist generates papers receiving higher review scores by DeepReviewer than existing fully automated systems. Nevertheless, we identify important limitations from the author evaluation and the Agents4Science reviews, indicating the potential risks of directly applying current AI Scientist systems and key challenges for future research. Finally, we comprehensively report various risks identified during development. We believe this study clarifies the current role and limitations of AI Scientist systems, offering insights into the areas that still require human expertise and the risks that may emerge as these systems evolve. Issues, comments, and questions are all welcome in https://github.com/Agent4Science-UTokyo/Jr.AI-Scientist.
@@ 1 Introduction
Understanding the current capabilities in AI Scientist systems (autoresearch), autonomous agents capable of conducting scientific research, is crucial for promoting sustainable, AI-driven scientific progress. Nevertheless, developers must remain conscious of the potential risks these systems pose to the academic ecosystem and commit to advancing them responsibly. Since 2025, a new venue dedicated to evaluating AI-driven scientific contributions, the Agents4Science conference (Zou et al., 2025; Bianchi et al., 2025), has emerged. Through such a platform, developers of AI Scientist systems are encouraged to engage in responsible research and development, ensuring both the protection of the academic ecosystem and the long-term sustainability of scientific progress.
@@
In recent years, several works have explored the concept of AI Scientists (Lu et al., 2024; Weng et al., 2025a; Yamada et al., 2025; Tang et al., 2025). However, the quality of research papers produced by these systems remains insufficient. One major reason is that the problem setting of achieving fully automated science is overly ambitious and often lacks clearly defined scientific goals for AI Scientists. Without a specific goal, these systems tend to generate undirected discoveries that appear to lack genuine scientific value. Another limitation is that current systems are limited to small-scale code experiments (Zhu et al., 2025b; Lu et al., 2024; Yamada et al., 2025), lacking the scale and complexity needed for meaningful science. Achieving real scientific contributions requires not just ideas but strong implementation capability to handle complex codebases.
@@
As an initial step toward enabling AI Scientists to produce genuine scientific value, we can take inspiration from how student researchers begin their research. When a student first joins a research lab, a common and meaningful process is as follows: the mentor assigns a key paper, the student analyzes its limitations, proposes an improvement hypothesis, implements the idea on the baseline code, iteratively experiments until improvements are achieved, and finally writes a paper summarizing the results. Through this process, the student learns the fundamental workflow of scientific research and gains the skills and experience needed for more creative work later on. Also, improving a baseline method is not only an important stage in early research training but also a valuable research goal in many fields where advancing task performance remains a central scientific pursuit.
@@
In this paper, we introduce Jr. AI Scientist, a new AI Scientist that mimics the essential research workflow of a novice student researcher: Starting from a baseline paper, it identifies key limitations, proposes an improvement hypothesis, iteratively conduct experiments until improvements are achieved, and writes a paper with the results. Specifically, A Jr. AI Scientist is defined as an AI Scientist that is given baseline resources and focuses on extending the baseline. This setting is novice student analog, where a student builds upon an existing paper which the mentor gave. The workflow of Jr. AI Scientist is shown in Figure˜1. We provide the baseline paper, its LaTeX source files, and the associated codebase for each stage. Table˜1 shows a comparison of the problem setting with existing research. This problem setting reframes previously ambitious goals into a more specific objective, providing a clear optimization direction for the AI Scientist. Moreover, because the framework operates on the actual codebase of baseline papers, it can generate results with genuine scientific value. These aspects collectively represent an essential first step toward the autonomous generation of reliable and high-quality research papers.
@@
Our Jr. AI Scientist consists of three main components: (1) automatic idea generation based on the limitations of a given paper, (2) automatic implementation and iteratively experiments until improvements are achieved, and (3) automatic writing of a research paper based on the obtained results. This system is built upon AI Scientist v2 (Yamada et al., 2025), but our work differs from prior studies (Lu et al., 2024; Weng et al., 2025a; Yamada et al., 2025; Tang et al., 2025) in several key aspects: First, by leveraging the latest coding agents (e.g. Claude Code (Anthropic, 2025a)), our system can perform meaningful improvements and edits on real multi-file codebases, which were challenging for previous AI Scientist systems. Second, by incorporating the full set of resources from a given baseline paper, the system exploits all available artifacts such as LaTeX sources, PDFs, and codebases, thereby substantially improving the scope and quality of every stage in the research pipeline. Finally, by refining every stage of the process, our framework enables the autonomous generation of research papers that are both higher in quality and more trustworthy.
@@
AI Scientist Systems Starting Point Code Complexity Review Score AI Scientist-v1 (Lu et al., 2024) Template code Single file 3.30 AI Scientist-v2 (Yamada et al., 2025) General idea Single file 2.75 AI Researcher (Tang et al., 2025) 15-20 existing papers Multiple files 3.25 Jr. AI Scientist (ours) One baseline paper and code Multiple files 5.75
@@
For the baseline papers, we selected papers for which we obtained permission from the original authors. Specifically, we used three papers: NeurIPS 2023 paper (Miyai et al., 2023) and IJCV 2025 paper (Miyai et al., 2025b) on out-of-distribution (OOD) detection (a task that aims to detect semantic classes outside the predefined set of semantic classes) (Hendrycks and Gimpel, 2017; Yang et al., 2024), and an ICLR 2025 spotlight paper (Zhang et al., 2025) on pre-training data detection for large language models (LLMs). Refer to Section˜4.1 for the detailed rationale behind the selection of these papers. For the evaluation, we conducted three evaluations: (1) an automated assessment using DeepReviewer (Zhu et al., 2025a), (2) an author-led evaluation, and (3) submission to the Agents4Science conference (Zou et al., 2025). DeepReviewer automatically compared our generated papers with existing AI-generated works to assess overall quality. The author evaluation examined the outputs for hallucinations or fabricated content. Finally, the Agents4Science (Zou et al., 2025) platform provides rigorous evaluation and feedback from the community platform.
@@
Through our experiments, the Jr. AI Scientist successfully generated new research papers that build upon the above top-tier venues works by proposing and implementing novel algorithms. As for the evaluation using DeepReviewer (Zhu et al., 2025a), the papers generated by Jr. AI Scientist achieved substantially higher review scores compared to the existing AI-generated papers. Therefore, our Jr. AI Scientist can be regarded as the most capable autonomous AI Scientist. However, we also observed that Jr. AI Scientist still exhibits some failures and unresolved challenges through the author evaluation and Agents4Science conference. To share these challenges and lessons with the research community, we analyze the feedback and evaluation results from the Agents4Science conference and include the author evaluation, which helps clarify what is required to further improve the quality of Jr. AI Scientist systems.
@@
Revealing Strengths and Limitations of Jr. AI Scientist: We conducted extensive evaluations using open-source AI reviewers, Agents4Science, and author evaluation. The results demonstrate that Jr. AI Scientist generates higher-quality research papers than existing AI Scientists, while also revealing key challenges for future improvement.
@@ 2 Related Work
Automated Scientific Discovery. Recent progress has significantly reshaped the role of AI in automating end-to-end scientific research. AI Scientist-v1 (Lu et al., 2024) was an early milestone, showcasing how advanced language models can autonomously generate research ideas, run experiments, and draft scientific papers. This work was followed by a series of subsequent studies in machine learning fields (Intology, 2025; Tang et al., 2025) and diverse scientific disciplines (Villaescusa-Navarro et al., 2025; Mitchener et al., 2025) that further advanced this line of research. However, these approaches often suffer from an overly ambitious problem setting that aims to achieve fully automated science and tend to lack clearly defined scientific objectives for AI Scientists. Without specific goals, such systems often produce undirected discoveries that lack genuine scientific value. To address this issue, our Jr. AI Scientist builds on existing baselines and conducts research within a well-defined research workflow, aiming to generate higher-quality scientific papers. As concurrent work, DeepScientist (Weng et al., 2025b) also adopts a baseline-based approach. DeepScientist (Weng et al., 2025b) focuses mainly on experimental performance, formalizing discovery as a Bayesian Optimization problem. However, its overall framework design and workflow integration are outlined at a high level, with limited discussion of implementation details. In contrast, our Jr. AI Scientist explicitly articulates each stage of the research process and further aims to contribute to the community by comprehensively reporting the failures and risks encountered throughout scientific exploration.
@@
AI-Assisted Scientific Research. Research specialized for each element of the research process has also been actively explored (Chen et al., 2025). For the idea generation phase, Si et al. (2025b) investigates the novelty of the LLM-generated ideas, and Si et al. (2025a) investigates the ideation–execution gap. For the survey phase, OpenScholar (Asai et al., 2026) have been developed to support literature review. For the experimental phase, AlphaEvolve (Novikov et al., 2025) leverages large-scale trial-and-error strategies to enhance the performance. For the writing and review phase, CycleResearcher (Weng et al., 2025a) provides a learning framework specialized for scientific writing, while DeepReviewer (Zhu et al., 2025a) focuses on the review process. Recent studies provide a comprehensive overview of automated review, outlining key challenges, proposing a practical review pipeline for real-world implementation, and constructing a large-scale dataset to support automated review research (Zhuang et al., 2025; Lin et al., 2023a; b). Beyond these, rather than pursuing full automation like AI Scientists, AI Co-Scientist (Gottweis et al., 2025) emphasizes collaboration between humans and AI. In this work, instead of focusing on individual parts of the research process, we investigate the entire end-to-end research cycle, aiming to rigorously evaluate both the performance and the limitations.
@@
Failures and Risk Analysis for AI Scientist Systems. There are very few studies that thoroughly analyze or report the risks and failure cases of AI Scientist systems. Although Tang et al. (2024) summarizes the risks that AI Scientists may pose, it focuses on hypothesis-based potential risks, rather than empirically observed risks or failures. While Beel et al. (2025) provides an in-depth analysis of failure cases in AI Scientist-v1 (Lu et al., 2024), these findings are based on the early AI Scientist, and the analysis was not conducted from the developer’s perspective. While Luo et al. (2025) examines four failure modes (benchmark selection, data leakage, metric misuse, and post-hoc bias), their analysis is limited to experimental diagnostics and early AI Scientists without modern coding. Thus, their analysis remains somewhat limited in scope.
@@ 3.1 Preparation: Baseline Paper Selection
The preparation stage involves selecting a baseline paper, obtaining its LaTeX source files and PDF, and the baseline code. This setup is realistic because many recent publications are released on arXiv with LaTeX sources, and their implementation code is shared on GitHub. While there might be some augment that AI agents should automatically select a baseline and reproduce the code, current reproducibility rates from papers are still limited (Siegel et al., 2025; Xiang et al., 2025; Starace et al., 2025), making complete automation premature. Since our goal is to emulate how a human scientist engages in early-stage research under the guidance of a mentor, we explicitly include this preparatory stage.
When constructing the baseline code, we followed AI Scientist v1 (Lu et al., 2024) and made only minor modifications to the existing implementation so that the experiments could be executed via baseline.py and the results could be visualized via plot.py. Defining such an experimental entry point facilitated easier management and reproducibility of the execution process in the experimental section.
@@ 3.2 Idea Generation Phase
We provide a baseline paper as text to an LLM (e.g. o4-mini (OpenAI, 2025b)) and prompt it to output the limitations of the work. Based on both the baseline paper and the limitations, the LLM is then guided to propose potential research ideas. Following AI Scientist v2 (Yamada et al., 2025), the system evaluates the originality of proposed ideas through literature review tools such as Semantic Scholar, which review papers citing the baseline work and papers with similar concepts. If conceptually similar ideas are identified, they are refined; otherwise, they are clearly distinguished from prior work. These steps define the preliminary research idea.
@@ 3.3.1 Preliminary: Coding Agent Usage
A powerful coding agent (e.g. Claude Code (Anthropic, 2025a)) is employed to translate research ideas into concrete implementations. We provide the coding agent with a working directory that contains the baseline implementations (prepared at Section˜3.1) and give it detailed instructions through input prompts. The agent is informed of how to use the main scripts—baseline.py, which serves as the experimental entry point, and plot.py, which visualizes the experimental results. We use claude-sonnet-4-20250514 within Claude Code (version 1.0.24).